OpenStudy (anonymous):

Let "dodeka" number be defined as a positive integer whose digits, when added together, equal 12. i) the number of dodeka even numbers between 10 and 100. ii) the number of dodeka odd numbers between 10 and 100. Which of these is true? A) choice i quantity is greater. B) choice ii quantity is greater. C) the two quantities are equal. D) the relationships can not be determined.

OpenStudy (anonymous):

I'm not sure how you'd mathematically solve this but, 1+9 = 10. so forget about #'s <=19 2+9 = 11, so forget about #'s<=29 3+9 = 12 so start at 39 i) has to be even so X(39) 48 CHECK X(57) 66 CHECK X(75) 84 CHECK X(93) , 9+9 = 18, 1+0+0 = 1 and I also showed " ii) " Pretty sure B is the answer. :)

OpenStudy (anonymous):

I just found the first #, 39, that adds to 12, then I just added 1 (to the tens) and subtracted 1 (from the ones) till I got to 100, or 94(or 9+4=13)(which is greater than 12), then figured out there weren't anymore after.

OpenStudy (anonymous):

got it thank you :D